# VLG-CBM: Training Concept Bottleneck Models with Vision-Language Guidance

### v1

@@ 4.2 A New Evaluation Metric for CBM: Number of Effective Concepts (NEC)
Eq. 8 provides a formal theoretical explanation on [18]’s observation. Moreover, it raises a concern on the evaluation of CBMs: model classification accuracy may not be a good metric for evaluating the semantical information learned in CBL, because a random CBL could also achieve high accuracy. To address this concern, we need to control the concept number $k$ so that the semantic meaningful CBLs can be distinguished with random CBLs w.r.t. the final classification performance.
@@
    [13, 20] suggested using a sparse linear layer for final prediction to enhance interpretability. Though sparsity is initially introduced to enhance interpretability, we note that this also reduces the number of concepts used in the decision, thus controlling the information leakage. The problem is, these works lack the quantification for sparsity, which is necessary for fair comparison between methods.
@@
Given these benefits, we suggest to control the NEC when comparing the performance of CBMs.
@@ 5.1 Performance comparison
Setup. Following prior work [13], we conduct experiments on five image recognition datasets: CIFAR10, CIFAR100[5], CUB[16], Places365[22] and ImageNet[15]. We follow the choice of [13] to use CLIP(RN50) [14] for CIFAR10 and CIFAR100, ResNet-18[2] for CUB and ResNet-50 (trained on ImageNet) for Places365 and ImageNet as the backbone.
@@
Results. The test results are summarized in Table 2 with the following observations:
@@
Dataset CIFAR10 CIFAR100 CUB200 Places365 ImageNet
Metrics Acc@5 Avg. Acc. Acc@5 Avg. Acc. Acc@5 Avg. Acc. Acc@5 Avg. Acc. Acc@5 Avg. Acc.
Random 67.55% 77.45% 29.52% 47.21% 68.91% 73.44% 17.57% 28.62% 41.49% 61.97%
LF-CBM 84.05% 85.43% 56.52% 62.24% 53.51% 69.11% 37.65% 42.10% 60.30% 67.92%
LM4CV 53.72% 69.02% 14.64% 36.70% N/A N/A N/A N/A N/A N/A
LaBo 78.69% 82.05% 44.82% 55.18% N/A N/A N/A N/A N/A N/A
VLG-CBM (Ours) 88.55% 88.63% 65.73% 66.48% 75.79% 75.82% 41.92% 42.55% 73.15% 73.98%
Table 2: Performance comparison between VLG-CBM, LF-CBM[13], LM4CV[18], LaBo[19] and a random baseline. The random baseline has 1024 neurons for CIFAR10 and CIFAR100, 512 for CUB, 2048 for Places365, and 4096 for ImageNet. The results of LM4CV and LaBo on CUB, Places365, and ImageNet are marked as "N/A" because they could not be applied on non-CLIP backbones.
@@ 5.2 Visualization of CBL neurons
In order to examine whether our CBL learns concepts aligned with human perception, we list the top-5 activated images for example concept neurons on the model trained on the CUB dataset in Fig. 4. As shown in the figure, our CBL faithfully captures the corresponding concept. We provide more visualization results in Appendix G.
@@ 5.3 Case study
In this section, we conduct a case study to compare the concept prediction between our VLG-CBM , LF-CBM [13] and LM4CV[18] as shown in Fig 1. We provide extended results and comparison with LaBo [19] in Section H.2. For our method, we use the final layer with NEC $=5$. We show that our method provides more accurate concept prediction and more interpretable decisions for users.
Decision interpretability. We examine the explanation of each CBM model on example images by showing the top-5 concept contributions. The contribution of each concept is calculated as the product of the concept prediction value and corresponding weight. Formally, the contribution of $i$-th concept to $j$-th class is defined as: $\text{Contribution}(i,j)=g_{i}(z)\cdot(W_{F})_{ji}$.
We pick the top-5 contributing concepts for the final predicted class and visualize it in Fig. 1. We could see that:
@@ Appendix D Evaluating annotations from Grounding DINO
This section quantitatively evaluates concept annotations obtained from Grounding DINO. We use CUB dataset for comparison which contains ground-truth for fine-grained concepts present in each image. We use the label set from Koh et al. [4] which has 1:1 mapping with the ground-truth concepts in the CUB dataset. We use precision and recall metric to measure the quality of annotations from Grouding DINO for each concepts. Table D.1 present mean precision and mean recall value at different confidence threshold. We observe that the obtained annotations have a very high recall i.e if the concept is present in the image, grounding DINO is able to retrieve the object. The precision is also sufficiently high though it suffers from a relatively higher false-positive detection rate compared to false-negative detection rate. However, as demonstrated in our qualitative and quantitative studies (Table 2, Fig 4, G.2, G.1) the effect of false-positive is minimal and VLG-CBMis able to faithfully represent concepts in the Concept Bottleneck Layer.
@@ Appendix F Constructing model with specified NEC
Appendix G Visualizing VLG-CBM explanations
This section presents an extended version of Table 4 visualizing top-5 images for randomly picked concepts for CUB and Places365 dataset. The results are shown in Fig G.1, G.3, G.4, and G.5 for Places365 and G.2, Fig G.6, Fig G.7, and Fig G.8 for CUB.
@@ Appendix G Visualizing VLG-CBM explanations
Figure G.1: Top-5 activating images for randomly selected Places365 concepts
@@
Figure G.2: Top-5 activating images for randomly selected CUB concepts
Figure G.3: Randomly selected explanations for Places365 (Part 1)
Figure G.4: Randomly selected explanations for Places365 (Part 2)
Figure G.5: Randomly selected explanations for Places365 (Part 3)
Figure G.6: Randomly selected explanations for CUB (Part 1)
Figure G.7: Randomly selected explanations for CUB (Part 2)
Figure G.8: Randomly selected explanations for CUB (Part 3)
Appendix H Additional case study examples
H.1 Negative concepts in reasoning
Figure H.1: Image 307: An example of negative reasoning of VLG-CBM
H.2 Impact of NEC
The study in Section 5.3 shows that our VLG-CBM provides more interpretable decisions than baseline methods. To better understanding where these advantages comes from, we conduct a further study to set the baselines with NEC=5 and compare the decision interpretation, see Figs. H.2, H.3 and H.4. The results suggest setting NEC=5 alleviate the problem from non-top-5 concepts. However, wrong/inaccurate/less useful explanations still exist.
Figure H.2: Comparing baselines with different NECs
Figure H.3: Comparing baselines with different NECs
Figure H.4: Comparing baselines with different NECs
Appendix I Further discussion on decision explanations
I.1 Negative contributions

### v2

@@ 4.2 A New Evaluation Metric for CBM: Number of Effective Concepts (NEC)
Eq. 8 provides a formal theoretical explanation on [25]’s observation. Moreover, it raises a concern on the evaluation of CBMs: model classification accuracy may not be a good metric for evaluating the semantical information learned in CBL, because a random CBL could also achieve high accuracy. To address this concern, we need to control the concept number $k$ so that the semantically meaningful CBLs can be distinguished with random CBLs w.r.t. the final classification performance.
@@
    [15, 28] suggested using a sparse linear layer for final prediction to enhance interpretability. Though sparsity is initially introduced to enhance interpretability, we note that this also reduces the number of concepts used in the decision, thus controlling the information leakage. However, the problem is, these works lack the quantification for sparsity, which is necessary for fair comparison between methods.
@@
Given these benefits, we suggest to control the NEC when comparing the performance of CBMs. In Section 5, we provide experiments with controlled NEC, where we observed our VLG-CBM  outperforms other baselines.
@@ 5.1 Performance comparison
Setup. Following prior work [15], we conduct experiments on five image recognition datasets: CIFAR10, CIFAR100[7], CUB[23], Places365[30] and ImageNet[18]. For the choice of backbone, we categorize the experiments into two categories:
  1. 1.
    CLIP backbone: For CLIP backbone, we choose CLIP-RN50 for all datasets.
  2. 2.
    non-CLIP backbobe: We follow the choice of [15] to use ResNet-18[3] for CUB and ResNet-50 (trained on ImageNet) for Places365 and ImageNet as the backbone.
The reason of this categorization is some previous works (e.g. LaBo[27] only supports CLIP backbone.
@@
Results. We summarize the test results for with backbone in Table 2 and results with non-CLIP backbones in Table 3. From the results, we could see that:
@@
Dataset CIFAR10 CIFAR100 ImageNet CUB
Metrics Acc@5 Avg. Acc Acc@5 Avg. Acc Acc@5 Avg. Acc Acc@5 Avg. Acc
Random 67.55% 77.45% 29.52% 47.21% 18.04% 39.63% 25.37% 40.13%
LF-CBM 84.05% 85.43% 56.52% 62.24% 52.88% 62.24% 31.35% 52.70%
LM4CV 53.72% 69.02% 14.64% 36.70% 3.77% 26.65% 3.63% 15.25%
LaBo 78.69% 82.05% 44.82% 55.18% 24.27% 45.53% 41.97% 59.27%
VLG-CBM(Ours) 88.55% 88.63% 65.73% 66.48% 59.74% 62.70% 60.38% 66.03%
Table 2: Performance comparison with CLIP-RN50 backbone. We compare our method with a random baseline, LF-CBM[15], LM4CV[25] and LaBo[27]. The random baseline has 1024 neurons for CIFAR10 and CIFAR100, 512 for CUB and 4096 for ImageNet.
Dataset CUB Places365 ImageNet
Metrics Acc@5 Avg. Acc. Acc@5 Avg. Acc. Acc@5 Avg. Acc.
Random 68.91% 73.44% 17.57% 28.62% 41.49% 61.97%
LF-CBM 53.51% 69.11% 37.65% 42.10% 60.30% 67.92%
VLG-CBM (Ours) 75.79% 75.82% 41.92% 42.55% 73.15% 73.98%
Table 3: Performance comparison with non-CLIP backbone. We compare against LF-CBM[15] and a random baseline, as LM4CV[25], LaBo[27] does not support non-CLIP backbone. The random baseline has 1024 neurons for CIFAR10 and CIFAR100, 512 for CUB, 2048 for Places365, and 4096 for ImageNet.
@@ 5.2 Visualization of CBL neurons
In order to examine whether our CBL learns concepts aligned with human perception, we list the top-5 activated images for example concept neurons on the model trained on the CUB dataset in Fig. 4. As shown in the figure, our CBL faithfully captures the corresponding concept. We provide more visualization results in Appendix K.
@@ 5.3 Case study
In this section, we conduct a case study to compare the concept prediction between our VLG-CBM , LF-CBM [15] and LM4CV[25] as shown in Fig 1. We provide extended results and comparison with LaBo [27] in Section G.2. For our method, we use the final layer with NEC $=5$. We show that our method provides more accurate concept prediction and more interpretable decisions for users.
Decision interpretability. We examine the explanation of each CBM model on example images by showing the top-5 concept contributions. The contribution of each concept is calculated as the product of the concept prediction value and corresponding weight. Formally, the contribution of $i$-th concept to $j$-th class is defined as: $\text{Contribution}(i,j)=g_{i}(z)\cdot(W_{F})_{ji}$. We pick the top-5 contributing concepts for the final predicted class and visualize it in Fig. 1. We could see that:
@@ 6 Conclusion, Potential Limitations and Future work
Acknowledgement
The authors thank the anonymous reviewers for valuable feedback on the manuscript. The authors are partially supported by National Science Foundation awards CCF-2107189, IIS-2313105, IIS-2430539, Hellman Fellowship, and Intel Rising Star Faculty Award. The authors also thank ACCESS for support in this work.
@@ Appendix D Evaluating annotations from Grounding DINO
This section quantitatively evaluates concept annotations obtained from Grounding DINO. We use CUB dataset for comparison which contains ground-truth for fine-grained concepts present in each image. We use the label set from Koh et al. [6] which has 1:1 mapping with the ground-truth concepts in the CUB dataset. We use precision and recall metric to measure the quality of annotations from Grouding DINO for each concepts. Table D.1 present mean precision and mean recall value at different confidence threshold. We observe that the obtained annotations have a very high recall i.e if the concept is present in the image, grounding DINO is able to retrieve the object. The precision is also sufficiently high though it suffers from a relatively higher false-positive detection rate compared to false-negative detection rate. However, as demonstrated in our qualitative and quantitative studies (Table 3, Fig 4, K.2, K.1) the effect of false-positive is minimal and VLG-CBMis able to faithfully represent concepts in the Concept Bottleneck Layer.
@@ Appendix F Constructing model with specified NEC
Appendix G Additional case study examples
G.1 Negative concepts in reasoning
Figure G.1: Image 307: An example of negative reasoning of VLG-CBM
G.2 Impact of NEC
The study in Section 5.3 shows that our VLG-CBM provides more interpretable decisions than baseline methods. To better understanding where these advantages comes from, we conduct a further study to set the baselines with NEC=5 and compare the decision interpretation, see Figs. G.2, G.3 and G.4. The results suggest setting NEC=5 alleviate the problem from non-top-5 concepts. However, wrong/inaccurate/less useful explanations still exist.
Figure G.2: Comparing baselines with different NECs
Figure G.3: Comparing baselines with different NECs
Figure G.4: Comparing baselines with different NECs
Appendix H Further discussion on decision explanations
H.1 Negative contributions
Appendix I Additional experiment results
I.1 Generalizability to OOD datasets
In this section, we study a question: will our VLG-CBM hurts the generalization ability of original model to Out-Of-Distribution(OOD) dataset? To study this problem, we conduct experiment on Waterbirds dataset [19]. Waterbirds is an OOD dataset adapted from the CUB dataset, which combines bird photos from CUB with image backgrounds from Places365. We use the same ResNet model as we used in Table 3 for CUB dataset. For VLG-CBM, we choose NEC=5 and compare the results with the standard, non-interpretable models. On this dataset, the results are shown below:
Method CUB Accuracy Waterbirds Accuracy
Standard model (black-box) 76.70% 69.83%
VLG-CBM 75.79% 69.83%
Table I.1: Accuracy of VLG-CBM and standard blackbox model on CUB and Waterbirds datasets.
It can be seen that our VLG-CBM generalizes well as the standard model does, which shows that our VLG-CBM is competitive and has very small accuracy trade-off with the interpretability compared with the standard black-box model.
I.2 Ablation study
I.2.1 Ablation on augmentation probability
In this section, we conduct an ablation study on the probability of applying our crop-to-concept data augmentation introduced in Section 3.1. The dataset we used is the CUB dataset and the backbone is ResNet as we used in the main experiment. The results are listed below.
Crop-to-Concept-Prob Acc@NEC=5 Avg. Acc
0.0 75.73 75.76
0.2 75.83 75.88
0.4 75.71 75.72
0.6 75.57 75.62
0.8 75.52 75.57
1.0 72.29 73.15
From the table, we could see that the performance is best with augmentation probability 0.2.
Appendix J Human study
In this section, we present a human study following the practice of Oikarinen et al. [15] on Amazon MTurk platform. To briefly summarize, we show the annotator top-5 contributing concepts of our method (VLG-CBM) and baseline (LF-CBM or LM4CV) and asking them which one is better. The scores for each method are assigned as 1-5 according to the response of annotators: 5 for the explanations from VLG-CBM is strongly more reasonable, 4 for VLG-CBM is slightly more reasonable, 3 for both models are equally reasonable, 2 for the baseline is slightly more reasonable, and 1 for the baseline is strongly more reasonable. Thus, if our model provides better explanations than the baselines, then we should see a score higher than 3. We show an example screenshot of our study in Fig. J.1.
We report the average score in Table J.1 for two baselines: LF-CBM and LM4CV. For each baseline, we randomly sample 200 images and collect 3 results from 3 different annotators. It can be seen that VLG-CBM has scores higher than 3 for both baselines, indicating our VLG-CBM provides better explanations than both baselines. LaBo is excluded in our experiment due to its dense layer and large number of concepts: the top-5 concepts usually account for less than 0.01% of final prediction.
Figure J.1: An example of human study interface
Experiment Score (VLG-CBM) Score (Baseline)
VLG-CBM vs. LF-CBM 3.33 (1.54) 2.67 (1.54)
VLG-CBM vs. LM4CV 3.38 (1.54) 2.62 (1.54)
Table J.1: Average Mturk score for our VLG-CBM and two baselines.
Appendix K Visualizing VLG-CBM explanations
This section presents an extended version of Table 4 visualizing top-5 images for randomly picked concepts for CUB and Places365 dataset. The results are shown in Fig K.1, K.3, K.4, and K.5 for Places365 and K.2, Fig K.6, Fig K.7, and Fig K.8 for CUB.
@@ I.2.1 Ablation on augmentation probability
Figure K.1: Top-5 activating images for randomly selected Places365 concepts
@@
Figure K.2: Top-5 activating images for randomly selected CUB concepts
Figure K.3: Randomly selected explanations for Places365 (Part 1)
Figure K.4: Randomly selected explanations for Places365 (Part 2)
Figure K.5: Randomly selected explanations for Places365 (Part 3)
Figure K.6: Randomly selected explanations for CUB (Part 1)
Figure K.7: Randomly selected explanations for CUB (Part 2)
Figure K.8: Randomly selected explanations for CUB (Part 3)
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our introduction in Section 1 summarizes the contribution and scope of this paper accurately.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide a full set of assumptions in our Eq. 8 and we provide complete and correct proof in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Section 5, we present the settings of our experiments. We present more implementation details on Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The link to the project webpage is provided in the abstract and the code will be released by poster deadline.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We include experimental details in Sections 5 and B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The experiments are very computationally expensive to repeat and measure the error bar.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide computational resources in Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research is conducted following NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification:This work focus on technical development of making neural network models more interpretable.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cite the dataset and models we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.